OpenStudy (anonymous):

★ WRITING Tell whether division is commutative and associative. Give examples to support your answer.

OpenStudy (anonymous):

Think about the commutative property. For multiplication you have 3*4=4*3. In division, would the same idea work? 3/4=4/3 And just like that, the associative property 3/(4/5) = (3/4)/5 What do you think, are the statemetns I've made true?